Evaluate
∫√(1+√x) dx
1
Expert's answer
2019-10-27T12:40:51-0400

Solution:


We need to find the Integration of given function.

We can use the suitable substitution to change to simple form


We have formulae about differentiation


"\frac {d} {dx} (x^n) = n \times x^ {n-1}\\
\frac {d}{dx} (constant ) =0"


"Let \space 1 + \sqrt x = t"

Differentiate with respect to x

"d(1 + \sqrt x) = dt"

"0 + \frac {1} {2 \sqrt x} dx = dt"

"\frac {1} {2 \sqrt x} dx = dt"

"dx = 2 \sqrt x dt"

"dx = 2 ( t -1 ) dt"

"\int \sqrt {( 1 + \sqrt x)} dx = \int (\sqrt t) 2 (t -1) dt"

"=2 \int ( t^ {\frac {3}{2} } -t^ {\frac {1}{2}} ) dt +c"

"= 2 ( \frac {t^ {\frac {3}{2} +1}}{\frac {3}{2} +1} - \frac {t^ {\frac {1}{2} +1}}{\frac {1}{2} +1} ) +c"


"= 2 ( \frac {t^ {\frac {5}{2} }}{\frac {5}{2}} - \frac {t^ {\frac {3}{2} }}{\frac {3}{2} } ) +c"

"= 4 ( \frac {t^ {\frac {5}{2} }}{5} - \frac {t^ {\frac {3}{2} }}{3} ) +c"




"= 4 ( \frac {(1+ \sqrt x)^ {\frac {5}{2} }}{5} - \frac {(1 + \sqrt x)^ {\frac {3}{2} }}{3} ) +c"

Answer :

"\int \sqrt {( 1 + \sqrt x)} dx = 4 ( \frac {(1+ \sqrt x)^ {\frac {5}{2} }}{5} - \frac {(1 + \sqrt x)^ {\frac {3}{2} }}{3} ) +c"